Circular No. 7866

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bz IN MCG +05-34-33 T. Puckett, Mountain Town, GA; and S. Gauthier, London, ON, re-

port the discovery of an apparent supernova (mag 16.9) on an unfiltered CCD frame (limiting mag 19.0) taken with the Puckett Observatory 0.35m automated supernova patrol telescope on Apr. 3.39 UT. SN 2002bz is located at α = 14^{h}24^{m}40^{s}.52, δ = +26^{o}37′34′′.7 (equinox 2000.0), which is 0′′.1 west and 3′′.7 north of the center of MCG +05-34-33. The new object was also present and confirmed on unfiltered CCD frames taken on Apr. 4.17, but it does not appear on unfiltered images taken by Puckett on Feb. 19 and Mar. 14 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1950 Apr. 12 (limiting mag ∼ 20.0) and on 1993 Mar. 30 and 1996 Apr. 19 (limiting mag ∼ 21.0).

SUPERNOVA 2002ca IN UGC 8521 Independent discoveries of an apparent supernova in UGC 8521 have

been reported from unfiltered CCD images taken by Puckett and B. Kerns, Stanton, ND, with the 0.35-m telescope (as above) on Apr. 3.29 and 4.41 UT (mag 17.3) and by W. D. Li, University of California at Berkeley, and LOTOSS (cf. IAUC 7514) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 3.3 and 4.3 (mag ∼ 17.3). Puckett reports the new object at α = 13^{h}32^{m}31^{s}.51, δ = +1^{o}50′42′′.5 (equinox 2000.0), which is 11′′.5 east and 9′′.8 south of the center of UGC 8521; Li provides position end figures 31^{s}.51, 41′′.7. Approximate limiting magnitudes for images showing nothing at this position: 1984 Mar. 7, 20.0 (Palomar Sky Survey, reported by Puckett); 1996 Mar. 25 and 1997 Apr. 8, 21.0 (ibid.); 2002 Feb. 17 and 23, 20.0 (Puckett, CCD); Mar. 5.3, 19.0 (KAIT).

SUPERNOVA 2002by IN ESO 139-G34 N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports

that spectra (range 380–540 nm) of SN 2002by (cf. IAUC 7865) were obtained by C. Gronwall and H. Ford (Johns Hopkins University) with the CTIO 1.5-m telescope (+ facility spectrograph) on Apr. 3.4 UT. The reduced spectrophotometry shows that this object is a supernova, but the small wavelength coverage and the weakness of the features do not allow a specific classification. The spectrum is possibly consistent with a type-I supernova, but redder spectra will be needed to give an definitive classification.

2002 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green